# Dynamical Formation of Apparent Horizons due to Boundary Effect in Vacuum Einstein Gravity

### v2

@@ 2. Geometric setup and semi-global development
It is convenient to decompose the entire Cauchy slice $\mathcal{M}_{t=-a}$ into three regions: an “interior” region $\widehat{\mathcal{M}}_{\mathrm{int}}$, lying to the past of $\mathcal{D}_{a,\epsilon}$; the interaction region $\widehat{\mathcal{M}}_{-a}$ defined above; and an “exterior” region $\widehat{\mathcal{M}}_{\mathrm{ext}}$, lying to the future/outside of $\mathcal{D}_{a,\epsilon}$. In particular, we write
(5) $\mathcal{M}_{t=-a}\;=\;\widehat{\mathcal{M}}_{\mathrm{int}}\;\cup\;\widehat{\mathcal{M}}_{-a}\;\cup\;\widehat{\mathcal{M}}_{\mathrm{ext}}.$
The decomposition (5) isolates the portion $\widehat{\mathcal{M}}_{-a}$ of the Cauchy hypersurface on which the large characteristic data propagate and interact, separating it from the complementary interior and exterior regions of $\mathcal{M}_{t=-a}$.
Figure 1. The schematics of the current framework: concentration of the generalized mean curvature $H-|\kappa|$ while increasing the radius. The initial data is provided on the null hypersurfaces $u=u_{\infty}$ and $\underline{u}=0$, and the interior Cauchy slice $\mathcal{M}_{1}=\widehat{\mathcal{M}}_{int}$. The Cauchy data on the slice $\widehat{\mathcal{M}}_{t=-a}$ is prescribed by gluing data on $\mathcal{M}_{1}$, the induced data on $\mathcal{M}_{2}$ by the Characteristic development on the slab $\mathcal{D}_{a,\epsilon}:=[u_{\infty},-a]\times[0,\epsilon]\times\mathbb{S}^{2}$, and exterior of a Kerr slice data on $\mathcal{M}_{3}$. The data on $\mathcal{M}_{1}$ is set such that it does not contain an apparent horizon. By the solution of the Characteristic semi-global problem in $\mathcal{D}_{a,\epsilon}$, one does not have any apparent horizon in the entire Cauchy slice $\widehat{\mathcal{M}}_{t=-a}$. In particular, $S_{-a,0}=\underline{H}_{0}\cap H_{-a}$ does not verify the condition of [38], while $S_{-a,\epsilon}:=H_{-a}\cap\underline{H}_{\epsilon}$ does. Therefore, $J^{+}(\mathcal{M}_{1})\cap\widehat{M}_{t=-a+\epsilon}$ has an apparent horizon inside.
3. Gluing of Cauchy data and horizon exclusion
@@
By the semi-global existence result established in the previous section, the development $\mathcal{D}_{a,\epsilon}$ induces canonical Cauchy data on the portion $\widehat{\mathcal{M}}_{-a}:=\mathcal{M}_{t=-a}\cap\mathcal{D}_{a,\epsilon}$. On the complementary regions $\widehat{\mathcal{M}}_{\mathrm{int}}$ and $\widehat{\mathcal{M}}_{\mathrm{ext}}$, one retains the freedom to prescribe data independently, subject only to the constraint and rigidity conditions imposed by the positive mass theorem [33, 34, 35]. In particular, the data on the exterior region $\widehat{\mathcal{M}}_{\mathrm{ext}}$ are far more constrained than those on the interior piece $\widehat{\mathcal{M}}_{\mathrm{int}}$, and a natural choice is to attach a Kerr exterior with prescribed ADM mass and angular momentum via the Corvino–Schoen gluing construction. The objective is then to glue the data on the three components
$\widehat{\mathcal{M}}_{\mathrm{int}},\qquad\widehat{\mathcal{M}}_{-a},\qquad\widehat{\mathcal{M}}_{\mathrm{ext}},$
@@
so as to produce a smooth global Cauchy data set on $\mathcal{M}_{t=-a}$ satisfying the Einstein constraint equations. Note that in the context of dynamical trapped surface formation [23] constructed an initial data set free of trapped surfaces for the Cauchy problem compatible with Christodoulou’s short pulse data type. Later, [10] constructed an initial data set for the dynamical formation of the apparent horizon (though this definition of apparent horizon is slightly different from ours in the sense that in their definition it is a spacelike hypersurface that is foliated by marginally outer trapped surfaces (i.e., $\mbox{tr}\chi=0$) whereas in our context, by an apparent horizon, we simply mean a spacelike topological 22-sphere for which $\mbox{tr}\chi=0,\mbox{tr}\underline{\chi}<0$) and subsequently addressed the Penrose inequality.
@@
The guiding idea is to choose the initial data on the null cones $H_{u_{\infty}}$, $\underline{H}_{0}$, and on the interior Cauchy region $\widehat{\mathcal{M}}_{\mathrm{int}}$ in such a way that the boundary of $\widehat{\mathcal{M}}_{\mathrm{int}}$ fails to satisfy the apparent–horizon condition of [38], namely,
(6) $c_{\partial\widehat{\mathcal{M}}_{\mathrm{int}}}:=\min_{\partial\widehat{\mathcal{M}}_{\mathrm{int}}}\big[H-|\kappa|\big]<\frac{3\pi}{2\,\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})}.$
The strict inequality (6) ensures that $\widehat{\mathcal{M}}_{\mathrm{int}}$ contains no apparent horizon. Assuming that throughout the semi–global region $\mathcal{D}_{a,\epsilon}$ one has
@@
and that the exterior Kerr portion $\widehat{\mathcal{M}}_{\mathrm{ext}}$ is also free of trapped surfaces, it follows that no apparent horizon occurs anywhere in the spacetime before or up to the slice t=at=-a.
Note that the semi–global domain $\mathcal{D}_{a,\epsilon}$ includes the causal past of the topological sphere $S_{-a,\epsilon}$. By transporting the geometric quantity $c$ along the incoming null direction, each sphere $S_{u,\underline{u}}$ near $(-a,\epsilon)$ can be focused so that $S_{-a,\epsilon}$ attains a large value of $c$. Concurrently, by the standard local existence theory for the Cauchy problem, one obtains a future development of the slice $\mathcal{M}_{t=-a}$, thereby generating a nearby Cauchy hypersurface $\mathcal{M}_{t=-a+\epsilon}$. For $\epsilon>0$ sufficiently small—chosen in a teleological manner depending on the prior data—the slice $\mathcal{M}_{t=-a+\epsilon}$ inherits analogous geometric estimates (e.g. curvature and lapse–shift bounds). Moreover, since $\mbox{tr}\chi>0$ along u=au=-a by the semi–global construction, the intersection of the causal future of $\widehat{\mathcal{M}}_{\mathrm{int}}$ with $\mathcal{M}_{t=-a+\epsilon}$ satisfies
(7) $\mathrm{Rad}\!\left(J^{+}(\widehat{\mathcal{M}}_{\mathrm{int}})\cap\mathcal{M}_{t=-a+\epsilon}\subset\mathcal{M}_{t=-a+\epsilon}\right)\approx\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})+C\epsilon,$
and for a constant $C$ dependent on the initial data and so for $\epsilon$ sufficiently small
(8) $\mathrm{Rad}\!\left(J^{+}(\widehat{\mathcal{M}}_{\mathrm{int}})\cap\mathcal{M}_{t=-a+\epsilon}\subset\mathcal{M}_{t=-a+\epsilon}\right)\approx\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})+\epsilon^{\frac{1}{2}},$
showing that the interior region expands monotonically towards the future for sufficiently small time. This is sufficient for our purpose.
We now analyze the geometric condition on the future Cauchy slice $\mathcal{M}_{t=-a+\epsilon}$ that signals the onset of trapped geometry. Observe that the boundary of the domain
$\widehat{\mathcal{M}}_{\mathrm{int}}\cap\mathcal{M}_{t=-a+\epsilon}$
within $\mathcal{M}_{t=-a+\epsilon}$ coincides with the topological two–sphere $S_{-a,\epsilon}$. Define, analogously to (6), the boundary curvature functional
$c_{S_{-a,\epsilon}}:=\min_{S_{-a,\epsilon}}\big[H-|\kappa|\big],$
where $H$ denotes the mean curvature of $S_{-a,\epsilon}$ in $\mathcal{M}_{t=-a+\epsilon}$ and $\kappa$ is the trace of the second fundamental form of $\mathcal{M}_{t=-a+\epsilon}$ restricted to $S_{-a,\epsilon}$. Suppose that
(9) $c_{S_{-a,\epsilon}}>\frac{3\pi\big(\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})-\epsilon^{\frac{1}{2}}\big)}{2\,\big[\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})\big]^{2}}.$
Then, according to the criterion of [38], the domain $\widehat{\mathcal{M}}_{\mathrm{int}}\cap\mathcal{M}_{t=-a+\epsilon}$ must contain an apparent horizon in its interior.
The key geometric mechanism underlying (9) is as follows. Rather than starting from a configuration dominated by large incoming radiation from past null infinity, we consider the complementary regime in which the outgoing radiation along the incoming null hypersurface $\underline{H}_{0}$ is dominant. Concretely, we assume that the squared shear $|\underline{\hat{\chi}}|^{2}$ is large on $\underline{H}_{0}$, which enhances focusing via the null structure equations. This strong shear, together with nonlinear couplings among the Ricci coefficients and curvature components, amplifies the quantity $H-|\kappa|$ at the boundary sphere $S_{-a,\epsilon}$. Simultaneously, as indicated by (6)–(8), The radius of the interior region increases from $\mathrm{Rad}(\widehat{\mathcal{M}}_{\mathrm{int}})$ to $\mathrm{Rad}\big(J^{+}(\widehat{\mathcal{M}}_{\mathrm{int}})\cap\mathcal{M}_{t=-a+\epsilon}\big)$. The combined effect is that strong outgoing radiation, concentrated within a suitably defined ringed region, distorts the boundary geometry of the interior domain, producing a surface with sufficiently large generalized mean curvature to satisfy the inequality (9). This mechanism provides a new geometric route to apparent-horizon formation.
@@
(16) $||\eta,\underline{\eta}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\omega||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{1}{|u|},~||\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{1}{|u|}.$
With this scaling hierarchy, we can integrate $\mbox{tr}\underline{\chi}$ along the incoming direction, i.e., work with the following equation
@@
which, upon integration and using the bounds 15 one obtains for every $\underline{u}\in[0,\epsilon]$ (provided one have the semi-global construction already)
(19) $\mbox{tr}\underline{\chi}(-a,\underline{u})=\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\underline{u})}{a}-\frac{1}{a}\int_{u_{\infty}}^{a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\underline{u})du^{{}^{\prime}}+\frac{C(\mathcal{I}^{0})}{a^{2}},$
where the constants involved depend only on the initial data. In particular, note that the term $\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}du^{{}^{\prime}}$ can make $\mbox{tr}\underline{\chi}(-a,\underline{u})$ large negative. Now we control $\mbox{tr}\chi$ through integrating along $\nabla_{4}$ direction. Recall that the Raichowdhury equation
@@
But the presence of $|\hat{\chi}|^{2}$ term with a negative sign can reduce $\mbox{tr}\chi$. As it turns out, controlling the size of $\epsilon$ and the new hierarchy where $\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim|u|^{-1}$, the reduction of $\mbox{tr}\chi$ can be controlled and this potentially dangerous term contributes to a negligible amount. Integration and using the bounds 15 yield
(21) $\mbox{tr}\chi(u,\epsilon)=\mbox{tr}\chi(u,0)-\int_{0}^{\epsilon}\bigg(-\frac{1}{2}(\mbox{tr}\chi)^{2}-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi\bigg)du^{{}^{\prime}}\approx\mbox{tr}\chi(u,0)-\frac{Ca\epsilon}{|u|^{2}}.$
Now we are at a place to actually compute $H-|\kappa|$ at $(-a,\epsilon)$ and (a,0)(-a,0). Using 11, we obtain
(22) $\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=\bigg(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\bigg)(-a,\epsilon)-\bigg|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigg|(-a,\epsilon)$
(23) $=\bigg(\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}-\frac{C}{a^{2}}\bigg)$
(24) $-\bigg|\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}+\frac{C}{a^{2}}\bigg|$

### v4

@@ 2. Overview and strategy of the proof
It is convenient to decompose the entire Cauchy slice $\mathcal{M}_{t=-a}$ into three regions: an “interior” region $\mathcal{M}_{1}$, lying in the interior; the interaction region $\mathcal{M}_{-a}$ defined above; and an “exterior” region $\mathcal{M}_{\mathrm{ext}}$, lying to the outside of $D_{a,1}$. In particular, we write
(6) $\mathcal{M}_{t=-a}\;=\;\mathcal{M}_{1}\;\cup\;\mathcal{M}_{2}\;\cup\;\mathcal{M}_{\mathrm{ext}}.$
The decomposition (6) isolates the portion $\mathcal{M}_{2}$ of the Cauchy hypersurface on which the large characteristic data propagate and interact, separating it from the complementary interior and exterior regions of $\mathcal{M}_{t=-a}$. One of the main parts of the current work is the proof of this semi-global development with a new hierarchy of data and concentration of the generalized Yau mean curvature along u=au=-a (or u=a1/au=-a-1/a) null cone. Now we recall the second ingredient in the construction of [44], the $H$-radius of Schoen-Yau radius $\text{Rad}(\Omega)$ of a domain $\Omega$.
Definition 3.
Let $\Gamma$ be a simply closed curve in $\Omega$ that bounds a disk in $\Omega$. We let $N_{r}(\Gamma)$ denote the set of points within a distance $r$ of $\Gamma$. We define the $H$-radius or Schoen-Yau radius of $\Omega$ with respect to $\Gamma$ as
$\text{Rad}(\Omega,\Gamma):=\sup\{r:\text{dist}(\Gamma,\partial\Omega)>r,~\Gamma\text{does not bound a disk in}N_{r}(\Gamma)\}.$
We define the Schoen-Yau radius or H-radius of $\Omega$ denoted by $\text{Rad}(\Omega)$ as follows
$\text{Rad}(\Omega):=\sup\{\text{Rad}(\Omega,\Gamma):\Gamma\text{as above}\}.$
The guiding principle behind the introduction of this radius is to quantify, in a geometrically intrinsic manner, the effective interior thickness of a domain $\Omega$, as opposed to its global diameter or volume. More precisely, the radius is designed to discriminate between domains that are uniformly thick in all directions and those that exhibit pronounced geometric anisotropy, such as long, thin, tube-like regions.
This distinction is already apparent in basic model geometries. If $\Omega\subset\mathbb{R}^{3}$ is a Euclidean ball of radius $R$, then the $H$-radius or Schoen-Yau radius satisfies
$\operatorname{Rad}(\Omega)=\frac{R}{2},$
reflecting the fact that the domain possesses comparable thickness in every direction. By contrast, if $\Omega$ is a cylindrical region of the form $\mathbb{S}^{2}_{R}\times(-L,L)$, where $\mathbb{S}^{2}_{R}$ denotes the round sphere of radius $R$, then a direct computation yields
$\operatorname{Rad}(\Omega)=\min\!\left(\frac{\pi R}{2},\,L\right).$
In particular, in the highly elongated regime $L\gg R$, the HH–radius is governed entirely by the transverse scale $R$, and is insensitive to the longitudinal extent of the domain. This behavior precisely captures the intended geometric content: the radius detects the maximal scale on which the domain remains uniformly thick, while deliberately ignoring directions along which the geometry degenerates into a thin tube.
Accordingly, the $H$-radius provides a robust quantitative measure of interior thickness that sharply distinguishes globally round geometries from elongated or neck–like configurations, a feature that will play a crucial role in the subsequent geometric and analytic arguments. With these geometric-topological notions, [44] proves a remarkable theorem on the existence of an MOTS on a Cauchy slice. More precisely, the theorem is as follows
Theorem 2.1.
[44] Let $M$ be a space-like hypersurface in a spacetime. Let $g_{ij}$ be its induced metric and $k_{ij}$ be its second fundamental form. Assume that the spacelike mean curvature $H$ of $\partial M$ is strictly greater than its time-like mean curvature $\mbox{tr}_{\partial M}k=\kappa$. Let $c:=\min_{\partial M}\bigg({H-|\kappa|}\bigg)$. If $c\geq\frac{3\pi}{2\text{Rad}(M)}$, then $M$ must admit a MOTS in its interior.
Note that this is purely a boundary effect in the sense that if the (generalized) boundary mean curvature of a large domain (in the sense of large Schoen-Yau radius) in the Cauchy slice is significantly higher, then an MOTS must exist inside. This leads to the following natural question: Can one start from a regular configuration free of any MOTSs and form an MOTS in an evolutionary manner in finite time. This statement needs to be made more precise, such as exactly what it means by a regular configuration, etc. We will do this momentarily.
Figure 2. A detailed diagram depicting the mechanism of producing a MOTS in an evolutionary manner: The schematics of the current framework: concentration of the generalized mean curvature $H-|\kappa|>0$ while increasing the radius. Consider the domain $\mathcal{M}^{1/a}_{int}\subset\mathcal{M}_{-a}$. The choice of initial data on this interior and the characteristic evolution are designed to produce quasi-isometry between $(\Phi_{T}(M^{1/a}_{int}),(\Phi^{-1}_{T})^{*}g)$ and $(M^{1/a}_{int},g)$ for T=a1/a+3/4T=-a-1/a+3/4 up to $\frac{g}{a^{\frac{3}{2}}}$ while the extra collar $\bigg(J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{-a-1/a+\epsilon}\bigg)\setminus\Phi(\mathcal{M}^{1/a}_{int})_{-a-1/a+\epsilon}$ in the causal future contributes an additional $O(1)$ thickness gain while the generalized mean curvature $H-|\kappa|$ is preserved up to $O(a^{-5/2})$ along u=a1/au=-a-1/a up to $\underline{u}=3/4$ triggering the sign flip of the entity $(H-|\kappa|)-\frac{3\pi}{2\text{Rad}(\Omega)}$ from $\mathcal{M}^{1/a}_{int}$ to $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{-a-1/a+3/4}$ and subsequent formation of a MOTS inside the later.
2.1. Yau [44] criterion for MOTS formation
@@
By the semi-global existence result established in the previous section, the development $D_{a,1}$ induces canonical Cauchy data on the portion $\mathcal{M}_{-a}:=\mathcal{M}_{t=-a}\cap D_{a,1}$. On the complementary regions $\mathcal{M}_{1}$ and $\mathcal{M}_{\mathrm{ext}}$, one retains the freedom to prescribe data independently, subject only to the constraint and rigidity conditions imposed by the positive mass theorem [39, 40, 41]. In particular, the data on the exterior region $\mathcal{M}_{\mathrm{ext}}$ are far more constrained than those on the interior piece $\mathcal{M}_{1}$, and a natural choice is to attach a Kerr exterior with prescribed ADM mass and angular momentum via the Corvino–Schoen gluing construction. The objective is then to glue the data on the three components
$\mathcal{M}_{\mathrm{1}},\qquad\mathcal{M}_{2},\qquad\mathcal{M}_{\mathrm{ext}},$
@@
so as to produce a smooth global Cauchy data set on $\mathcal{M}_{t=-a}$ satisfying the Einstein constraint equations. Note that in the context of dynamical trapped surface formation, [28] first constructed an initial data set free of trapped surfaces for the Cauchy problem compatible with Christodoulou’s short pulse data type. Later, [11] constructed an initial data set for the dynamical formation of the MOTS and addressing a spacetime Penrose inequality.
@@
The guiding idea is to choose the initial data on the null cones $H_{u_{\infty}}$, $\underline{H}_{0}$, and on the interior Cauchy region $\mathcal{M}_{1}$ in such a way that the boundary of $\mathcal{M}_{1}$ fails to satisfy the MOTS condition of [44], namely,
(7) $c_{\partial\mathcal{M}_{1}}:=\min_{\partial\mathcal{M}_{1}}\big[H-|\kappa|\big]<\frac{3\pi}{2\,\mathrm{Rad}(\mathcal{M}_{1})}.$
The strict inequality (7) ensures that $\mathcal{M}_{1}$ contains no MOTS. This prescription is perfectly compatible with Causality since the data in the interior $\widetilde{\mathcal{M}}_{1}$ is not affected by that on $H_{u_{\infty}}$ and $\underline{H}_{0}$. After having proven that throughout the semi–global region $D_{a,1}$ one has
@@
and the fact that the exterior Kerr portion $\mathcal{M}_{\mathrm{ext}}$ is also free of closed trapped surfaces, it follows that no MOTS occurs anywhere in the spacetime before or up to the slice t=at=-a.
The vital point to note here is that it is rather difficult to simultaneously control the generalized mean curvature $c$ and the radius $\text{Rad}(M)$ such that $c\cdot\text{Rad}(M)\geq\frac{3\pi}{2}$. This is precisely why we consider the two-step process. Let us revert our attention back to the diagram 2. The initial data is provided on the hypersurfaces $\underline{u}=0$ (or $\underline{H}_{0}$), $u=u_{\infty}$ (or $H_{u_{\infty}}$), and $\mathcal{M}_{1}$ (or $\widetilde{\mathcal{M}}_{1}$). The strategy that we adopt here is that first, we want to control the generalized mean curvature $c$ along u=au=-a. In the next step, we will control the radius of slice $\mathcal{M}_{1}$ (or in pactice we will consider a modified domain $\mathcal{M}^{1/a}_{int}$) as it evolves in the future. Notice that along u=au=-a, evolution from $s_{-a,0}$ to $S_{-a,\epsilon}$ indicates moving in the outgoing direction and hence naively expect a non-concentration of the generalized mean curvature $c$. The first idea is whether one can preserve the generalized mean curvature along u=au=-a up to error terms $O(a^{-5/2})$. Then, naturally, one can impose the initial condition on $\mathcal{M}_{1}$ in such a way that 7 holds, but since along u=au=-a, the radius of the bounding Cauchy slices increases, the condition as stated in the theorem 2.1 can be satisfied for $J^{+}(\mathcal{M}^{1/a}_{int})\cap\mathcal{M}_{t=-a-1/a+\epsilon}$. Therefore, the two main challenges in this study are the proof of the semi-global development with the choice of a new hierarchy of initial data and controlling the Cauchy evolution upto time $a+\epsilon$ for $\epsilon=O(1)$. We illustrate these steps next.
2.2. Focusing $H-|\mbox{tr}_{S}k|$
@@
(14) $||\eta,\underline{\eta}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\omega||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{\frac{1}{2}}}{|u|^{2}},~||\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim\frac{a^{-\frac{1}{2}}}{|u|}.$
and the Weyl curvature scaling
(15) $|\alpha|=O(a^{-\frac{1}{2}}|u|^{-1}),~|\beta|=O(|u|^{-2}),~|\underline{\beta}|=O(a|u|^{-4}),~|(\rho,\sigma)|=O(a^{\frac{1}{2}}|u|^{-3}),~|\underline{\alpha}|=O(a^{\frac{3}{2}}|u|^{-5})$
Here, note importantly that we will have to prove that the involved constants in $\lesssim$ depend only on the initial scale-invariant data of the respective Ricci coefficients, and this hierarchy is propagated in the domain of spacetime constructed via semi-global development. With this scaling hierarchy, we can integrate $\mbox{tr}\underline{\chi}$ along the incoming direction, i.e., work with the following equation
@@
which, upon integration and using the bounds (13) one obtains for every $\underline{u}\in[0,\epsilon]$ (provided one have the semi-global construction already)
(18) $\mbox{tr}\underline{\chi}(-a,\underline{u})=\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\underline{u})}{a}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\underline{u})du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\bigg(\mbox{tr}\underline{\chi}+\frac{2}{|u^{{}^{\prime}}|}\bigg)du^{{}^{\prime}}+\mathscr{E},$
where the error term $\mathscr{E}$ is $O(a^{-\frac{5}{2}})$. Now we need to control $\widetilde{\mbox{tr}\underline{\chi}}:=\mbox{tr}\underline{\chi}+\frac{2}{|u|}$. Recall the following equation verified by $\widetilde{\mbox{tr}\underline{\chi}}$
$\nabla_{3}\widetilde{\mbox{tr}\underline{\chi}}+\mbox{tr}\underline{\chi}\widetilde{\mbox{tr}\underline{\chi}}=\frac{2}{|u|^{2}}(\Omega^{-1}-1)+|\widetilde{\mbox{tr}\underline{\chi}}|^{2}+2\underline{\omega}\mbox{tr}\underline{\chi}-|\underline{\hat{\chi}}|^{2}$
and subsequently
$\nabla_{3}(|u|^{2}\widetilde{\mbox{tr}\underline{\chi}})=2(\Omega^{-1}-1)+2|u|^{2}\underline{\omega}\mbox{tr}\underline{\chi}-|u|^{2}|\underline{\hat{\chi}}|^{2}$
which yields
(19) $|u|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u,\underline{u})=|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\underline{u})-\int_{u_{\infty}}^{u}|u^{{}^{\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime}}+O(\frac{a^{\frac{1}{2}}}{|u|}).$
Substitute (19) into (18) and obtain
(20) $\mbox{tr}\underline{\chi}(-a,\underline{u})=\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\underline{u})}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\underline{u})}{a^{2}}-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\underline{u})du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E},$
where the error term $\mathscr{E}=O(a^{-\frac{5}{2}})$ and negligible for sufficiently large $a\gg 1$. Now we control $\mbox{tr}\chi$ through integrating along $\nabla_{4}$ direction. Recall that the Raichoudhury equation
@@
But the presence of $|\hat{\chi}|^{2}$ term with a negative sign can reduce $\mbox{tr}\chi$. As it turns out, controlling the size of $\epsilon$ and the new hierarchy where $\hat{\chi}||_{L^{\infty}(S_{u,\underline{u}})}\lesssim a^{-\frac{1}{2}}|u|^{-1}$, the reduction of $\mbox{tr}\chi$ can be controlled and this potentially dangerous term contributes to a negligible amount. Integration and using the bounds 13 yield for $0<\epsilon\leq 1$
(22) $\mbox{tr}\chi(u,\epsilon)=\mbox{tr}\chi(u,0)+\int_{0}^{\epsilon}\bigg(-\frac{1}{2}(\mbox{tr}\chi)^{2}-|\hat{\chi}|^{2}-2\omega\mbox{tr}\chi\bigg)du^{{}^{\prime}}\approx\mbox{tr}\chi(u,0)-\frac{4\epsilon}{|u|^{2}}.$
Here $\approx$ means up to terms that decay strictly faster than $|u|^{-2}$. Note that these terms can be made much smaller than the $O(|u|^{-2})$ terms and therefore can be ignored at a heuristic level. Now we are at a place to actually compute $H-|\kappa|$ at $(-a,\epsilon)$ and (a,0)(-a,0). Using (9) and absorbing $\frac{1}{2}$ to the left, we obtain
(23) $2\bigg(H-|\mbox{tr}_{\Sigma}K|\bigg)(-a,\epsilon)=\bigg(\mbox{tr}\chi-\mbox{tr}\underline{\chi}\bigg)(-a,\epsilon)-\bigg|\mbox{tr}\chi+\mbox{tr}\underline{\chi}\bigg|(-a,\epsilon)$
$=\bigg(\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}-\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}-\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\epsilon)}{a^{2}}$
$+\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}+\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg)$
$-\bigg|\mbox{tr}\chi(-a,0)-\frac{C\epsilon}{a^{2}}+\frac{|u_{\infty}|\mbox{tr}\underline{\chi}(u_{\infty},\epsilon)}{a}+\frac{|u_{\infty}|^{2}\widetilde{\mbox{tr}\underline{\chi}}(u_{\infty},\epsilon)}{a^{2}}$
$-\frac{1}{a}\int_{u_{\infty}}^{-a}|u^{{}^{\prime}}||\underline{\hat{\chi}}|^{2}(u^{{}^{\prime}},\epsilon)du^{{}^{\prime}}-\frac{1}{a}\int_{u_{\infty}}^{-a}\frac{1}{|u^{{}^{\prime}}|^{2}}\int_{u_{\infty}}^{u^{{}^{\prime}}}|u^{{}^{\prime\prime}}|^{2}|\underline{\hat{\chi}}|^{2}du^{{}^{\prime\prime}}du^{{}^{\prime}}+\mathscr{E}\bigg|$